### v1

# Intersections of sumsets in additive number theory

Abstract.
Let AA be a subset of an additive abelian semigroup and let hAhA be the $h$-fold sumset of AA. The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in the semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. When does one have
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for some or all $h\geq 2$?
Key words and phrases:
Sumset, Haar measure of sumsets, additive number theory, combinatorial number theory
2000 Mathematics Subject Classification:
11B13, 11B05, 11B75, 11P70, 22D99
Supported in part by PSC-CUNY Research Award Program grant 66197-00 54.
1. The sumset intersection problem
Let $\mathbf{N}$, $\mathbf{N}_{0}$, and $\mathbf{Z}$ denote, respectively, the sets of positive integers, nonnegative integers, and integers.
A sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$. A decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ is strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$ and asymptotically strictly decreasing if the sequence is not eventually constant, that is, if $A_{q}\neq A_{q+1}$ for infinitely many $q$.
Let AA be a subset of an additive abelian semigroup. The $h$-fold sumset of AA is the set
$hA=\{a_{1}+\cdots+a_{h}:a_{i}\in A\text{ for all }i=1,\ldots,h\}.$
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets in the semigroup and let
$A=\bigcap_{q=1}^{\infty}A_{q}.$
For all $q\in\mathbf{N}$, we have $A\subseteq A_{q}$ and so $hA\subseteq hA_{q}$. It follows that
(1) $hA=h\left(\bigcap_{q=1}^{\infty}A_{q}\right)\subseteq\bigcap_{q=1}^{\infty}hA_{q}.$
This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, if $q\in\mathbf{N}$ and
$A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}$
then
$n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\prime}$
and
n=(nq)+(h2)(q)+((h1)q)-n=(-n-q)+(h-2)(-q)+((h-1)q)
for all $n\in\mathbf{N}_{0}$, and so $hA^{\sharp}_{q}=\mathbf{Z}$ for all $h\geq 2$. Let AA be any finite set of integers. The sumset hAhA is finite for all $h\in\mathbf{N}$. The asymptotically strictly decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ defined by
$A_{q}=A\cup A^{\sharp}_{q}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}$
satisfies $A=\bigcap_{q=1}^{\infty}A_{q}$. We have $hA_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$ and so
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
We want to know when equality replaces inclusion in relation (1).
Problem 1.
Describe, classify, understand the strictly decreasing sequences $(A_{q})_{q=1}^{\infty}$ of subsets of an additive abelian semigroup such that
(2) $h\left(\bigcap_{q=1}^{\infty}A_{q}\right)=\bigcap_{q=1}^{\infty}hA_{q}.$
For related work, see [1][11].
2. Intersections of sets of integers
Theorem 1.
Let AA be a set of nonnegative integers. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
We must show that if $n\in\bigcap_{q=1}^{\infty}hA_{q}$, then $n\in hA$.
The number of representations of $n$ as a sum of $h$ nonnegative integers is finite. Indeed, the set $\mathcal{R}(n)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in\mathbf{N}_{0}^{h}$ such that
$n=b_{1}+\cdots+b_{h}$
satisfies $|\mathcal{R}(n)|=\binom{n+h-1}{h-1}$. For all $q\in\mathbf{N}$, we have $n\in hA_{q}$ and so there exists $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}(n)$ such that
$n=a_{1,q}+\cdots+a_{h,q}.$
Because $\mathcal{R}(n)$ is finite, there is some $h$-tuple $(b_{1},\ldots,b_{h})\in\mathcal{R}(n)$ such that
$(b_{1},\ldots,b_{h})=(a_{1,q},\ldots,a_{h,q})$
for infinitely many $q$ and so $b_{i}=a_{i,q}\in A_{q}$ for infinitely many $q$. Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $q$ and so $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $n=b_{1}+\cdots+b_{h}\in hA$. This completes the proof. ∎
The set AA of integers is a basis of order $h$ for $\mathbf{Z}$ if $hA=\mathbf{Z}$. The set AA is a basis for $\mathbf{Z}$ if AA is a basis of order $h$ for $\mathbf{Z}$ for some positive integer $h$. If AA is a basis of order $h$ for $\mathbf{Z}$ and if $A\subseteq A^{\prime}$, then $A^{\prime}$ is a basis of order $h^{\prime}$ for $\mathbf{Z}$ for all $h^{\prime}\geq h$.
Theorem 2.
Let AA be a set of integers.
  1. (1)
    If AA is a basis of order $h_{0}$ for $\mathbf{Z}$ and if $(A_{q})_{q=1}^{\infty}$ is any decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then
    $hA=\bigcap_{q=1}^{\infty}hA_{q}$
    for all $h\geq h_{0}$.
  2. (2)
    If AA is not a basis of order $h_{0}$ for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
    $h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
  3. (3)
    If AA is not a basis for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
    $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
    for all $h\in\mathbf{N}$.
Proof.
(1) Let AA be a basis of order $h_{0}$ for $\mathbf{Z}$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For all $h\geq h_{0}$ and for all $q\in\mathbf{N}$, we have $hA=hA_{q}=\mathbf{Z}$ and so
$hA=\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
(2) If AA is not a basis of order $h_{0}$ for $\mathbf{Z}$, then $h_{0}A\neq\mathbf{Z}$. For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. It follows that
$h_{0}A\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
(3) If AA is not a basis for $\mathbf{Z}$, then $hA\neq\mathbf{Z}$ for all $h\in\mathbf{N}$. For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. It follows that
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\in\mathbf{N}$. This completes the proof. ∎
Theorem 3.
Let AA be an infinite set of integers that is bounded below and that does not contain all sufficiently large integers. There exists an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$.
Proof.
By Theorem 1, we can assume that $a^{*}=\min(A)<0$.
For all $q\in\mathbf{N}$, let
$A^{\flat}_{q}=\{r\in\mathbf{N}:r\geq q\}$
and let
$A_{q}=A\cup A^{\flat}_{q}=A\cup\{r\in\mathbf{N}:r\geq q\}.$
The sequence $(A_{q})_{q=1}^{\infty}$ is decreasing and $A=\bigcap_{q=1}^{\infty}A_{q}$.
If $q\geq 1$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$. Because AA does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, we have
$hA\subseteq hA_{q}\subseteq hA\cup\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}$
and
$\bigcap_{q=1}^{\infty}\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}=\emptyset.$
Therefore,
$hA\subseteq\bigcap_{q=1}^{\infty}hA_{q}\subseteq hA\cup\bigcap_{q=1}^{\infty}\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}=hA.$
This completes the proof. ∎
Problem 2.
Let AA be an infinite set of integers that is not a basis of order $h$ for $\mathbf{Z}$. Does there exist a strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$?
Problem 3.
For what sets AA of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
Problem 4.
For what sets AA of integers does there exist an asymptotically decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
3. Sumsets in locally compact groups
Theorem 4.
Let AA be a nonempty subset of a locally compact abelian group $G$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of closed compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for every positive integer $h$.
Proof.
For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, the sumset $hA_{q}$ and the sumset intersection $\bigcap_{q=1}^{\infty}hA_{q}$ are nonempty closed compact sets. Let $x\in\bigcap_{q=1}^{\infty}hA_{q}$. We must show that $x\in hA$.
For all $q\in\mathbf{N}$, we have $x\in hA_{q}$ and so there is an $h$-tuple
$(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\subseteq A_{1}^{h}$
such that
$x=a_{1,q}+\cdots+a_{h,q}.$
Consider the sequences $(a_{i,q})_{q=1}^{\infty}$ for $i=1,\ldots,h$. Because the set $A_{1}$ is compact and $a_{i,q}\in A_{1}$ for all $i=1,\ldots,h$ and $q\in\mathbf{N}$, there is an infinite strictly increasing sequence of positive integers $(q_{j})_{j=1}^{\infty}$ such that the $h$ subsequences $(a_{i,q_{j}})_{j=1}^{\infty}$ converge:
$\lim_{j\rightarrow\infty}a_{i,q_{j}}=x_{i}$
for all $i=1,\ldots,h$. Then
$x$ $=\lim_{j\rightarrow\infty}\left(a_{1,q_{j}}+\cdots+a_{h,q_{j}}\right)$
$=\left(\lim_{j\rightarrow\infty}a_{1,q_{j}}\right)+\cdots+\left(\lim_{j\rightarrow\infty}a_{h,q_{j}}\right)$
$=x_{1}+\cdots+x_{h}.$
Suppose that $x_{i}\notin A$ for some ii. Let $q_{0}=0$. Construct the sequence $\left(x_{i,q}\right)_{q=1}^{\infty}$ as follows;
$x_{i,q}=a_{i,q_{j}}\qquad\text{for $q\in[q_{j-1}+1,q_{j}]$.}$
Because $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence, we have $x_{i,q}\in A_{q}$ for all $q\in\mathbf{N}$ and
$\lim_{i\rightarrow\infty}x_{i,q}=x_{i}.$
The group $G$ is locally compact. Because AA is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from AA such that $x_{i,q}\in K$ for all sufficiently large $q$. Then $x_{i,q}\in K\cap A_{q}$ for all sufficiently large $q$. It follows that, for all $Q\in\mathbf{N}$, there exists $q_{1}\geq Q$ with $x_{i,q_{1}}\in K\cap A_{q_{1}}\subseteq K\cap A_{q}$ for all $q\in[1,Q]$, and so
$x_{i,q_{1}}\in\bigcap_{q=1}^{Q}\left(K\cap A_{q}\right)\neq\emptyset.$
Because the sets $K\cap A_{q}$ are compact, the finite intersection property implies
$\emptyset\neq\bigcap_{q=1}^{\infty}\left(K\cap A_{q}\right)=K\cap\left(\bigcap_{q=1}^{\infty}A_{q}\right)=K\cap A=\emptyset$
which is absurd. Therefore, $x_{i}\in A$ for all $i=1,\ldots,h$ and $h\in hA$. This completes the proof. ∎
Theorem 5.
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of nonempty compact sets in $\mathbf{R}^{n}$ and let $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
This follows from Theorem 4 with $G=\mathbf{R}^{n}$. ∎
Theorem 6.
Let $G$ be a locally compact abelian group with Haar measure $\mu$. For $h\in\mathbf{N}$, let $\theta_{h}\geq 0$ and let $(\theta_{h,q})_{q=1}^{\infty}$ be a sequence of nonnegative real numbers such that $\lim_{q\rightarrow\infty}\theta_{h,q}=\theta_{h}$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of compact subsets of $G$ such that $\mu(hA_{q})=\theta_{h,q}$ for all $q\in\mathbf{N}$. Let $A=\bigcap_{q=1}^{\infty}A_{q}$. Then $\mu(hA)=\theta_{h}$.
Proof.
For all $h\in\mathbf{N}$, the sequence $(hA_{q})_{q=1}^{\infty}$ is a decreasing sequence of compact subsets of $G$. By Theorem 4,
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
and so
$\mu(hA)=\mu\left(\bigcap_{q=1}^{\infty}hA_{q}\right)=\lim_{q\rightarrow\infty}\mu(hA_{q})=\lim_{q\rightarrow\infty}\mu(\theta_{h,q})=\theta_{h}.$
This completes the proof. ∎
4. Semigroup problems
Problem 5.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. Describe the set $\{h\in\mathbf{N}:hA=\bigcap_{q=1}^{\infty}hA_{q}\}$.
Problem 6.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h+1)A=\bigcap_{q=1}^{\infty}(h+1)A_{q}$?
When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h-1)A=\bigcap_{q=1}^{\infty}(h-1)A_{q}$?

### v2

@@
 Abstract.
-Let AA be a subset of an additive abelian semigroup and let hAhA be the $h$-fold sumset of AA.
+Let $A$ be a subset of an additive abelian semigroup and let $hA$ be the $h$-fold sumset of $A$.
 The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in the semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
@@
 Let $\mathbf{N}$, $\mathbf{N}_{0}$, and $\mathbf{Z}$ denote, respectively, the sets of positive integers, nonnegative integers, and integers.
-A sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$.
-A decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ is strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$ and asymptotically strictly decreasing if the sequence is not eventually constant, that is, if $A_{q}\neq A_{q+1}$ for infinitely many $q$.
-Let AA be a subset of an additive abelian semigroup. The $h$-fold sumset of AA is the set
+Let $S$ be an additive abelian semigroup and let $A$ be a subset of $S$.
+For every positive integer $h$, the $h$-fold sumset of $A$ is the set of all sums of $h$ not necessarily distinct elements of $A$, that is,
 $hA=\{a_{1}+\cdots+a_{h}:a_{i}\in A\text{ for all }i=1,\ldots,h\}.$
-Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets in the semigroup and let
+For all $x\in S$, we denote by $r_{A,h}(x)$ the number of representations of $x$ as a sum of $h$ elements of $A$, that is,
+$r_{A,h}(x)=\operatorname{\text{card}}\left\{(a_{1},\ldots,a_{h})\in A^{h}:a_{1}+\cdots+a_{h}=x\right\}.$
+The subset $A$ is a basis of order $h$ for $S$ if $hA=S$ (equivalently, if $r_{A,h}(x)\geq 1$ for all $x\in S$), and a basis for $S$ if $A$ is a basis of order $h$ for $S$ for some positive integer $h$.
+The subset $A$ is a nonbasis of order $h$ for $S$ if $A$ is not a basis of order $h$ for $S$, that is, if $hA\neq S$.
+The set $A$ is a nonbasis for $S$ if $A$ is a nonbasis of order $h$ for $S$ for every positive integer $h$.
+The sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$, and strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$.
+The sequence $(A_{q})_{q=1}^{\infty}$ asymptotically strictly decreasing if the sequence is not eventually constant, that is, if $A_{q}\neq A_{q+1}$ for infinitely many $q$.
+Every asymptotically strictly decreasing sequence contains a strictly decreasing subsequence.
+Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets in the semigroup $S$ and let
 $A=\bigcap_{q=1}^{\infty}A_{q}.$
@@
 (1) $hA=h\left(\bigcap_{q=1}^{\infty}A_{q}\right)\subseteq\bigcap_{q=1}^{\infty}hA_{q}.$
-This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, if $q\in\mathbf{N}$ and
-$A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}$
-then
-$n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\prime}$
+This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, for every positive integer $q$, we define
+$A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}.$
+Let $h\geq 2$. For every nonnegative integer $n$, we have
+$n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\sharp}_{q}$
 and
-n=(nq)+(h2)(q)+((h1)q)-n=(-n-q)+(h-2)(-q)+((h-1)q)
-for all $n\in\mathbf{N}_{0}$, and so $hA^{\sharp}_{q}=\mathbf{Z}$ for all $h\geq 2$.
-Let AA be any finite set of integers.
-The sumset hAhA is finite for all $h\in\mathbf{N}$.
+$n=(-n-q)+(h-2)(-q)+((h-1)q)\in hA^{\sharp}_{q}$
+and so $hA^{\sharp}_{q}=\mathbf{Z}$.
+Let $A$ be any finite set of integers.
+The sumset $hA$ is finite for all $h\in\mathbf{N}$ and so $hA\neq\mathbf{Z}$.
 The asymptotically strictly decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ defined by
 $A_{q}=A\cup A^{\sharp}_{q}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}$
-satisfies $A=\bigcap_{q=1}^{\infty}A_{q}$. We have $hA_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$ and so
+satisfies $A=\bigcap_{q=1}^{\infty}A_{q}$. Because $hA_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$ , we have
 $hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
@@
 Theorem 1.
-Let AA be a set of nonnegative integers.
+Let $A$ be a set of nonnegative integers.
 Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
@@
 Proof.
-We must show that if $n\in\bigcap_{q=1}^{\infty}hA_{q}$, then $n\in hA$.
-The number of representations of $n$ as a sum of $h$ nonnegative integers is finite. Indeed, the set $\mathcal{R}(n)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in\mathbf{N}_{0}^{h}$ such that
+We must show that for all $h\geq 2$, if $n\in\bigcap_{q=1}^{\infty}hA_{q}$, then $n\in hA$.
+The number $r_{\mathbf{N}_{0},h}(n)$ of representations of $n$ as a sum of $h$ nonnegative integers is finite.
+Indeed, the set $\mathcal{R}(n)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in\mathbf{N}_{0}^{h}$ such that
 $n=b_{1}+\cdots+b_{h}$
-satisfies $|\mathcal{R}(n)|=\binom{n+h-1}{h-1}$. For all $q\in\mathbf{N}$, we have $n\in hA_{q}$ and so there exists $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}(n)$ such that
+satisfies $r_{\mathbf{N}_{0},h}(n)=|\mathcal{R}(n)|=\binom{n+h-1}{h-1}$.
+For all $q\in\mathbf{N}$, we have $n\in hA_{q}$ and so there is an $h$-tuple $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}(n)$ such that
 $n=a_{1,q}+\cdots+a_{h,q}.$
@@
 $(b_{1},\ldots,b_{h})=(a_{1,q},\ldots,a_{h,q})$
-for infinitely many $q$ and so $b_{i}=a_{i,q}\in A_{q}$ for infinitely many $q$.
-Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $q$ and so $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $n=b_{1}+\cdots+b_{h}\in hA$.
+for infinitely many $q$ and so, for infinitely many $q$, we have $b_{i}=a_{i,q}\in A_{q}$ for all $i=1,\ldots,h$.
+Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $i=1,\ldots,h$ and for all $q$.
+It follows that $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $n=b_{1}+\cdots+b_{h}\in hA$.
 This completes the proof.
 ∎
-The set AA of integers is a basis of order $h$ for $\mathbf{Z}$ if $hA=\mathbf{Z}$.
-The set AA is a basis for $\mathbf{Z}$ if AA is a basis of order $h$ for $\mathbf{Z}$ for some positive integer $h$.
-If AA is a basis of order $h$ for $\mathbf{Z}$ and if $A\subseteq A^{\prime}$, then $A^{\prime}$ is a basis of order $h^{\prime}$ for $\mathbf{Z}$ for all $h^{\prime}\geq h$.
 Theorem 2.
-Let AA be a set of integers.
-  1. (1)
-    If AA is a basis of order $h_{0}$ for $\mathbf{Z}$ and if $(A_{q})_{q=1}^{\infty}$ is any decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then
+Let $S$ be an additive abelian semigroup such that $r_{S,h}(x)$ is finite for all $h\in\mathbf{N}$ and $x\in S$.
+Let $A$ be a subset of $S$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
+For every positive integer $h$,
+$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
+Proof.
+It suffices to observe that the proof of Theorem 1 uses only the finiteness of the representation function $R_{\mathbf{N}_{0},h}(n)$ for all $h$ and $n$. This result follows immediately. ∎
+Theorem 3.
+Let $S$ be an additive abelian semigroup and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
+If $A$ is a basis of order $h_{0}$ for $S$, then
 $hA=\bigcap_{q=1}^{\infty}hA_{q}$
 for all $h\geq h_{0}$.
-  2. (2)
-    If AA is not a basis of order $h_{0}$ for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
+Proof.
+For all $h\geq h_{0}$ and for all $q\in\mathbf{N}$, we have $hA=hA_{q}=S$. and so
+$hA=\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
+This completes the proof. ∎
+Theorem 4.
+Let $A$ be a set of integers.
+If $A$ is a nonbasis of order $h_{0}$ for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
 $h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
-  3. (3)
-    If AA is not a basis for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
+If $A$ is a nonbasis for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
 $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
@@
 Proof.
-(1) Let AA be a basis of order $h_{0}$ for $\mathbf{Z}$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
-For all $h\geq h_{0}$ and for all $q\in\mathbf{N}$, we have $hA=hA_{q}=\mathbf{Z}$ and so
-$hA=\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
-(2) If AA is not a basis of order $h_{0}$ for $\mathbf{Z}$, then $h_{0}A\neq\mathbf{Z}$. For all $q\in\mathbf{N}$, let
-$A_{q}=A\cup A^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
-Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. It follows that
+For all $q\in\mathbf{N}$, let
+$A_{q}=A\cup A_{q}^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
+Then $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
+Moreover, $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$.
+If $A$ is a nonbasis of order $h_{0}$ for $\mathbf{Z}$, then
 $h_{0}A\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
-(3) If AA is not a basis for $\mathbf{Z}$, then $hA\neq\mathbf{Z}$ for all $h\in\mathbf{N}$. For all $q\in\mathbf{N}$, let
-$A_{q}=A\cup A^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
-Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. It follows that
-$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}$
+If $A$ is a nonbasis for $\mathbf{Z}$, then for all $h\in\mathbf{N}$ we have
+$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
+This completes the proof. ∎
+Theorem 4 extends to the group of lattice points $\mathbf{Z}^{d}$ and to a more general class of semigroups. For the group $\mathbf{Z}^{d}$, we let
+$A_{q}=\left\{(n_{1},\ldots,n_{d})\in\mathbf{Z}^{d}:|n_{i}|\geq q\text{ for all }q\in\mathbf{N}\right\}.$
+Theorem 5.
+Let $S$ be an additive abelian semigroup such that $S$ contains a strictly decreasing sequence of sets $\left(A_{q}^{\sharp}\right)_{q=1}^{\infty}$ such that $\bigcap_{q=1}^{\infty}A_{q}^{\sharp}=\emptyset$ and $hA_{q}^{\sharp}=S$ for all $h\geq 2$ and $q\geq 1$.
+If $A$ is a nonbasis of order $h_{0}$ for $S$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
+$h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
+If $A$ is a nonbasis for $S$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
+$hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
+for all $h\in S$.
+Proof.
+If $A$ is a nonbasis of order $h_{0}$ for $S$, then $h_{0}A\neq S$. For all $q\in\mathbf{N}$, let
+$A_{q}=A\cup A_{q}^{\sharp}.$
+Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=S$ for all $q\in\mathbf{N}$. It follows that
+$h_{0}A\neq S=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
+If $A$ is a nonbasis for $S$, then for all $h\in\mathbf{N}$ we have $hA\neq S$ and
+$hA\neq S=\bigcap_{q=1}^{\infty}hA_{q}$
 for all $h\in\mathbf{N}$. This completes the proof. ∎
 Theorem 6.
-Let AA be an infinite set of integers that is bounded below and that does not contain all sufficiently large integers.
+Let $A$ be an infinite set of integers that is bounded below and that does not contain all sufficiently large integers.
 There exists an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$.
@@
 If $q\geq 1$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$.
-Because AA does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
+Because $A$ does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
 For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, we have
@@
 Problem 2.
-Let AA be an infinite set of integers that is not a basis of order $h$ for $\mathbf{Z}$.
+Let $A$ be an infinite set of integers that is a nonbasis of order $h$ for $\mathbf{Z}$.
 Does there exist a strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$?
 Problem 3.
-For what sets AA of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
+For what sets $A$ of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
 Problem 4.
-For what sets AA of integers does there exist an asymptotically decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
+For what sets $A$ of integers does there exist an asymptotically decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
 3. Sumsets in locally compact groups
 Theorem 7.
-Let AA be a nonempty subset of a locally compact abelian group $G$.
+Let $A$ be a nonempty subset of a locally compact abelian group $G$.
 Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of closed compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
@@
 $=x_{1}+\cdots+x_{h}.$
-Suppose that $x_{i}\notin A$ for some ii. Let $q_{0}=0$. Construct the sequence $\left(x_{i,q}\right)_{q=1}^{\infty}$ as follows;
+Suppose that $x_{i}\notin A$ for some $i$. Let $q_{0}=0$. Construct the sequence $\left(x_{i,q}\right)_{q=1}^{\infty}$ as follows;
 $x_{i,q}=a_{i,q_{j}}\qquad\text{for $q\in[q_{j-1}+1,q_{j}]$.}$
@@
 The group $G$ is locally compact.
-Because AA is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from AA such that $x_{i,q}\in K$ for all sufficiently large $q$.
+Because $A$ is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from $A$ such that $x_{i,q}\in K$ for all sufficiently large $q$.
 Then $x_{i,q}\in K\cap A_{q}$ for all sufficiently large $q$.
@@
 When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h-1)A=\bigcap_{q=1}^{\infty}(h-1)A_{q}$?
+I thank Diego Marques for the observation that led to Theorem 2.

### v3

@@
 Abstract.
-Let $A$ be a subset of an additive abelian semigroup and let $hA$ be the $h$-fold sumset of $A$.
-The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in the semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
+Let $A$ be a subset of an additive abelian semigroup $S$ and let $hA$ be the $h$-fold sumset of $A$.
+The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in $S$ and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
 When does one have
@@
 Key words and phrases:
-Sumset, Haar measure of sumsets, additive number theory, combinatorial number theory
+Sumsets in semigroups, Haar measure of sumsets, additive number theory, combinatorial number theory
 2000 Mathematics Subject Classification:
@@
 $r_{A,h}(x)=\operatorname{\text{card}}\left\{(a_{1},\ldots,a_{h})\in A^{h}:a_{1}+\cdots+a_{h}=x\right\}.$
-The subset $A$ is a basis of order $h$ for $S$ if $hA=S$ (equivalently, if $r_{A,h}(x)\geq 1$ for all $x\in S$), and a basis for $S$ if $A$ is a basis of order $h$ for $S$ for some positive integer $h$.
+The subset $A$ is a basis of order $h$ for $S$ if $hA=S$, or, equivalently, if $r_{A,h}(x)\geq 1$ for all $x\in S$.
+The subset $A$ is a basis for $S$ if $A$ is a basis of order $h$ for $S$ for some positive integer $h$.
 The subset $A$ is a nonbasis of order $h$ for $S$ if $A$ is not a basis of order $h$ for $S$, that is, if $hA\neq S$.
-The set $A$ is a nonbasis for $S$ if $A$ is a nonbasis of order $h$ for $S$ for every positive integer $h$.
-The sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$, and strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$.
-The sequence $(A_{q})_{q=1}^{\infty}$ asymptotically strictly decreasing if the sequence is not eventually constant, that is, if $A_{q}\neq A_{q+1}$ for infinitely many $q$.
+The set $A$ is a nonbasis for $S$ if $hA\neq S$ for every positive integer $h$.
+The sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$.
+The decreasing sequence $(A_{q})_{q=1}^{\infty}$ is strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$ and asymptotically strictly decreasing if $A_{q}\neq A_{q+1}$ for infinitely many $q$.
+A decreasing sequence is asymptotically strictly decreasing if and only if it is not eventually constant.
 Every asymptotically strictly decreasing sequence contains a strictly decreasing subsequence.
-Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets in the semigroup $S$ and let
+Let $(A_{q})_{q=1}^{\infty}$ be an asymptotically strictly decreasing sequence of sets in the semigroup $S$ and let
 $A=\bigcap_{q=1}^{\infty}A_{q}.$
@@
 This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, for every positive integer $q$, we define
-$A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}.$
-Let $h\geq 2$. For every nonnegative integer $n$, we have
+(2) $A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}.$
+Let $h\geq 2$. For every nonnegative integer $n$, we have $\pm(n+q)\in A^{\sharp}_{q}$ and $\pm(h-1)q\in A^{\sharp}_{q}$. It follows that
 $n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\sharp}_{q}$
 and
-$n=(-n-q)+(h-2)(-q)+((h-1)q)\in hA^{\sharp}_{q}$
+$n=-(n+q)+(h-2)(-q)+((h-1)q)\in hA^{\sharp}_{q}$
 and so $hA^{\sharp}_{q}=\mathbf{Z}$.
@@
 $hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
+Problem 1.
 We want to know when equality replaces inclusion in relation (1).
-Problem 1.
-Describe, classify, understand the strictly decreasing sequences $(A_{q})_{q=1}^{\infty}$ of subsets of an additive abelian semigroup such that
+The basic problem is to describe, classify, understand the strictly decreasing sequences $(A_{q})_{q=1}^{\infty}$ of subsets of an additive abelian semigroup such that
 (3) $h\left(\bigcap_{q=1}^{\infty}A_{q}\right)=\bigcap_{q=1}^{\infty}hA_{q}.$
 For related work, see [2][13].
-2. Intersections of sets of integers
+2. Intersections of sumsets in discrete groups and semigroups
 Theorem 1.
-Let $A$ be a set of nonnegative integers.
-Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
+Let $h\in\mathbf{N}$.
+Let $S$ be an additive abelian semigroup such that $r_{S,h}(x)$ is finite for all $x\in S$.
+Let $A$ be a subset of $S$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
+Then
+$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
+Proof.
+For all $x\in S$, the set $\mathcal{R}_{S,h}(x)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in S^{h}$ such that
+$x=b_{1}+\cdots+b_{h}$
+is finite with cardinality $|\mathcal{R}_{S,h}(x)|=r_{S,h}(x)$.
+Let $x\in\bigcap_{q=1}^{\infty}hA_{q}$.
+We must show that $x\in hA$.
+For all $q\in\mathbf{N}$, we have $x\in hA_{q}$ and so there is an $h$-tuple $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}_{S,h}(x)$ such that
+$x=a_{1,q}+\cdots+a_{h,q}.$
+Because $\mathcal{R}_{S,h}(x)$ is a finite set, there is some $h$-tuple $(b_{1},\ldots,b_{h})\in\mathcal{R}_{S,h}(x)$ such that
+$(b_{1},\ldots,b_{h})=(a_{1,q},\ldots,a_{h,q})$
+for infinitely many $q$ and so, for infinitely many $q$, we have $b_{i}=a_{i,q}\in A_{q}$ for all $i=1,\ldots,h$.
+Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $i=1,\ldots,h$ and $q\in\mathbf{N}$.
+It follows that $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $x=b_{1}+\cdots+b_{h}\in hA$.
+This completes the proof.
+∎
+Let $\mathbf{Z}^{d}$ be the additive abelian group of $d$-dimensional lattice points in $\mathbf{R}^{d}$ and let $\mathbf{N}_{0}^{d}$ be the additive abelian semigroup of $d$-dimensional nonnegative lattice points.
+Theorem 2.
+Let $d\geq 1$ and let $A$ be a subset of $\mathbf{N}_{0}^{d}$.
+Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $\mathbf{N}_{0}^{d}$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
 For every positive integer $h$,
@@
 Proof.
-We must show that for all $h\geq 2$, if $n\in\bigcap_{q=1}^{\infty}hA_{q}$, then $n\in hA$.
-The number $r_{\mathbf{N}_{0},h}(n)$ of representations of $n$ as a sum of $h$ nonnegative integers is finite.
-Indeed, the set $\mathcal{R}(n)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in\mathbf{N}_{0}^{h}$ such that
-$n=b_{1}+\cdots+b_{h}$
-satisfies $r_{\mathbf{N}_{0},h}(n)=|\mathcal{R}(n)|=\binom{n+h-1}{h-1}$.
-For all $q\in\mathbf{N}$, we have $n\in hA_{q}$ and so there is an $h$-tuple $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}(n)$ such that
-$n=a_{1,q}+\cdots+a_{h,q}.$
-Because $\mathcal{R}(n)$ is finite, there is some $h$-tuple $(b_{1},\ldots,b_{h})\in\mathcal{R}(n)$ such that
-$(b_{1},\ldots,b_{h})=(a_{1,q},\ldots,a_{h,q})$
-for infinitely many $q$ and so, for infinitely many $q$, we have $b_{i}=a_{i,q}\in A_{q}$ for all $i=1,\ldots,h$.
-Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $i=1,\ldots,h$ and for all $q$.
-It follows that $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $n=b_{1}+\cdots+b_{h}\in hA$.
-This completes the proof.
-∎
-Theorem 2.
-Let $S$ be an additive abelian semigroup such that $r_{S,h}(x)$ is finite for all $h\in\mathbf{N}$ and $x\in S$.
-Let $A$ be a subset of $S$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
-For every positive integer $h$,
-$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
-Proof.
-It suffices to observe that the proof of Theorem 1 uses only the finiteness of the representation function $R_{\mathbf{N}_{0},h}(n)$ for all $h$ and $n$. This result follows immediately. ∎
+For all $n\in\mathbf{N}_{0}$, the number $r_{\mathbf{N}_{0},h}(n)$ of representations of $n$ as a sum of $h$ nonnegative integers is finite.
+Indeed, $r_{\mathbf{N}_{0},h}(n)=\binom{n+h-1}{h-1}$.
+For all $\mathbf{n}=(n_{1},\ldots,n_{d})\in\mathbf{N}_{0}^{d}$, we have
+$r_{\mathbf{N}_{0}^{d},h}(\mathbf{n})=\prod_{i=1}^{d}r_{\mathbf{N}_{o},h}(n_{i})<\infty.$
+The result follows from Theorem 1. ∎
 Theorem 3.
@@
 If $A$ is a basis of order $h_{0}$ for $S$, then
+$h_{0}A=\bigcap_{q=1}^{\infty}h_{0}A_{q}$
+If the subgroup $S$ contains an additive identity and if $A$ is a basis of order $h_{0}$ for $S$, then
 $hA=\bigcap_{q=1}^{\infty}hA_{q}$
@@
 Proof.
-For all $h\geq h_{0}$ and for all $q\in\mathbf{N}$, we have $hA=hA_{q}=S$. and so
-$hA=\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
-This completes the proof. ∎
+For all $q\in\mathbf{N}$, we have $A\subseteq A_{q}$ and so $S=h_{0}A\subseteq h_{0}A_{q}\subseteq S$. It follows that $h_{0}A_{q}=S$ and
+$h_{0}A=S=\bigcap_{q=1}^{\infty}hA_{q}.$
+Let $h\geq h_{0}$.
+If $S$ contains an additive identity, then $h_{0}A\subseteq hA$.
+If $A$ is a basis of order $h_{0}$ for $S$, then $S=h_{0}A\subseteq hA\subseteq S$ and so $A$ is a basis of order $h$ for $S$.
+This completes the proof.
+∎
+Note that the additive semigroup $\mathbf{N}$ of positive integers does not contain an additive identity.
+The set $A=\mathbf{N}$ is a basis of order 1 for $\mathbf{N}$ but not a basis of order $h$ for any $h\geq 2$.
 Theorem 4.
-Let $A$ be a set of integers.
-If $A$ is a nonbasis of order $h_{0}$ for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
-$h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
-If $A$ is a nonbasis for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
+Let $d\in\mathbf{N}$ and let $G$ be any subgroup of $\mathbf{R}^{d}$ that contains $\mathbf{Z}^{d}$.
+Let $A$ be a subset of $G$.
+If $A$ is a nonbasis of order $h_{0}$ for $G$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and
+$h_{0}A\neq\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
+If $A$ is a nonbasis for $G$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and
 $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
-for all $h\in\mathbf{N}$.
+for all $h\geq 2$.
 Proof.
+For all $\mathbf{x}=(x_{1},\ldots,x_{d})\in\mathbf{R}^{d}$, define $\|\mathbf{x}\|=\max(|x_{i}|:i=1,\ldots,d)$. For all $q\in\mathbf{N}$, let
+$A_{q,d}^{\sharp}=\{\mathbf{x}\in G:\|\mathbf{x}\|\geq q\}.$
+Note that $\{\mathbf{x}\in\mathbf{Z}^{d}:\|\mathbf{x}\|\geq q\}\subseteq A_{q,d}^{\sharp}$.
+Let $\mathbf{x}=(x_{1},\ldots,x_{d})\in G$. If $x_{1}\geq 0$, then
+$\mathbf{y}=(x_{1}+(h-1)q,x_{2},\ldots,x_{d})\in A_{q}^{\sharp}$
+and
+$\mathbf{z}=(-q,0,\ldots,0)\in A_{q}^{\sharp}$
+and so
+$\mathbf{x}$ $=(x_{1}+(h-1)q,x_{2},\ldots,x_{d})+(h-1)(-q,0,\ldots,0)$
+$=\mathbf{y}+(h-1)\mathbf{z}\in hA_{q,d}^{\sharp}.$
+If $x_{1}\leq 0$, then
+$(x_{1}-(h-1)q,x_{2},\ldots,x_{d})\in A_{q}^{\sharp}$
+and
+$(q,0,\ldots,0)\in A_{q}^{\sharp}$
+and so
+$\mathbf{x}$ $=(x_{1}-(h-1)q,x_{2},\ldots,x_{d})+(h-1)(q,0,\ldots,0)$
+$=\mathbf{y}+(h-1)\mathbf{z}\in hA_{q,d}^{\sharp}.$
+Thus, $hA_{q,d}^{\sharp}=G$ for all $q\in\mathbf{N}$ and $h\geq 2$.
 For all $q\in\mathbf{N}$, let
-$A_{q}=A\cup A_{q}^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
-Then $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
-Moreover, $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$.
-If $A$ is a nonbasis of order $h_{0}$ for $\mathbf{Z}$, then
-$h_{0}A\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
-If $A$ is a nonbasis for $\mathbf{Z}$, then for all $h\in\mathbf{N}$ we have
-$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
-This completes the proof. ∎
-Theorem 4 extends to the group of lattice points $\mathbf{Z}^{d}$ and to a more general class of semigroups. For the group $\mathbf{Z}^{d}$, we let
-$A_{q}=\left\{(n_{1},\ldots,n_{d})\in\mathbf{Z}^{d}:|n_{i}|\geq q\text{ for all }q\in\mathbf{N}\right\}.$
-Theorem 5.
-Let $S$ be an additive abelian semigroup such that $S$ contains a strictly decreasing sequence of sets $\left(A_{q}^{\sharp}\right)_{q=1}^{\infty}$ such that $\bigcap_{q=1}^{\infty}A_{q}^{\sharp}=\emptyset$ and $hA_{q}^{\sharp}=S$ for all $h\geq 2$ and $q\geq 1$.
-If $A$ is a nonbasis of order $h_{0}$ for $S$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
-$h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
-If $A$ is a nonbasis for $S$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
-$hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
-for all $h\in S$.
-Proof.
-If $A$ is a nonbasis of order $h_{0}$ for $S$, then $h_{0}A\neq S$. For all $q\in\mathbf{N}$, let
-$A_{q}=A\cup A_{q}^{\sharp}.$
-Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=S$ for all $q\in\mathbf{N}$. It follows that
-$h_{0}A\neq S=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
-If $A$ is a nonbasis for $S$, then for all $h\in\mathbf{N}$ we have $hA\neq S$ and
-$hA\neq S=\bigcap_{q=1}^{\infty}hA_{q}$
-for all $h\in\mathbf{N}$. This completes the proof. ∎
+$A_{q}=A\cup A_{q,d}^{\sharp}.$
+Then $hA_{q}=hA_{q,d}^{\sharp}=G$ for all $q\in\mathbf{N}$ and $h\geq 2$.
+The sequence $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence of sets such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
+If $A$ is a nonbasis of order $h_{0}$ for $G$, then
+$h_{0}A\neq G=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
+If $A$ is a nonbasis for $G$, then
+$hA\neq G=\bigcap_{q=1}^{\infty}hA_{q}$
+for all $h\geq 2$. This completes the proof. ∎
 Theorem 5.
@@
 Proof.
-By Theorem 1, we can assume that $a^{*}=\min(A)<0$.
+Note that a set of integers that is bounded below cannot be a basis for $\mathbf{Z}$. By Theorem 2 with $d=1$, we can assume that $a^{*}=\min(A)<0$.
 For all $q\in\mathbf{N}$, let
@@
 The sequence $(A_{q})_{q=1}^{\infty}$ is decreasing and $A=\bigcap_{q=1}^{\infty}A_{q}$.
-If $q\geq 1$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$.
+If $q\in\mathbf{N}$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$.
 Because $A$ does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
@@
 This completes the proof. ∎
+Let $S$ be an additive abelian semigroup.
+A subset $A$ of $S$ is a maximal nonbasis of order $h$ for $S$ if $A$ is not a basis of order $h$ for $S$ but every proper superset of $A$ is a basis of order $h$ for $S$.
+For example, for every integer $h\geq 2$, the set $h\ast\mathbf{Z}=\{h\ell:\ell\in\mathbf{Z}\}$ of multiples of $h$ is a maximal nonbasis of order $h$ for $\mathbf{Z}$ if and only if $h$ is prime.
+Maximal nonbases were introduced by Nathanson [4] and the first nontrivial examples were constructed by Nathanson [4] and Erdős and Nathanson [1].
+Theorem 6.
+Let $A$ be a subset of the infinite abelian semigroup $S$.
+If $A$ is a maximal nonbasis of order $h$ for $S$ and if $(A_{q})_{q=1}^{\infty}$ is an asymptotically strictly decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$.
+Proof.
+Let $A$ be a maximal nonbasis of order $h$ for $S$ and let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
+For all $q$, the set $A$ is a proper subset of $A_{q}$ and so $hA_{q}=S$.
+It follows that $hA\neq S=\bigcap_{q=1}^{\infty}hA_{q}$.
+This completes the proof.
+∎
 Problem 2.
-Let $A$ be an infinite set of integers that is a nonbasis of order $h$ for $\mathbf{Z}$.
-Does there exist a strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$?
-Problem 3.
-For what sets $A$ of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
-Problem 4.
-For what sets $A$ of integers does there exist an asymptotically decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
-3. Sumsets in locally compact groups
+Let $A$ be a subset of the infinite abelian semigroup $S$ such that $A$ is a nonbasis of order $h$ but not a maximal nonbasis of order $h$ for $S$.
+Does there exist an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$?
+In particular, consider this problem for $S=\mathbf{Z}$.
+3. Intersections of sumsets in locally compact groups
 Theorem 7.
-Let $A$ be a nonempty subset of a locally compact abelian group $G$.
-Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of closed compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$.
-Then
+Let $A$ be a nonempty subset of a locally compact abelian group $G$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
 $hA=\bigcap_{q=1}^{\infty}hA_{q}$
@@
 The group $G$ is locally compact.
-Because $A$ is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from $A$ such that $x_{i,q}\in K$ for all sufficiently large $q$.
+Because $A=\bigcap_{q=1}^{\infty}A_{q}$ is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from $A$ such that $x_{i,q}\in K$ for all sufficiently large $q$.
 Then $x_{i,q}\in K\cap A_{q}$ for all sufficiently large $q$.
@@
 $\emptyset\neq\bigcap_{q=1}^{\infty}\left(K\cap A_{q}\right)=K\cap\left(\bigcap_{q=1}^{\infty}A_{q}\right)=K\cap A=\emptyset$
-which is absurd. Therefore, $x_{i}\in A$ for all $i=1,\ldots,h$ and $h\in hA$. This completes the proof. ∎
+which is absurd. Therefore, $x_{i}\in A$ for all $i=1,\ldots,h$ and $x\in hA$. This completes the proof. ∎
 Theorem 8.
-Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of nonempty compact sets in $\mathbf{R}^{n}$ and let $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
+Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of nonempty compact sets in $\mathbf{R}^{d}$ such that $A=\bigcap_{q=1}^{\infty}A_{q}\neq\emptyset$. For every positive integer $h$,
 $hA=\bigcap_{q=1}^{\infty}hA_{q}.$
 Proof.
-This follows from Theorem 7 with $G=\mathbf{R}^{n}$. ∎
+This follows from Theorem 7 with $G=\mathbf{R}^{d}$. ∎
 Theorem 9.
 Let $G$ be a locally compact abelian group with Haar measure $\mu$.
-For $h\in\mathbf{N}$, let $\theta_{h}\geq 0$ and let $(\theta_{h,q})_{q=1}^{\infty}$ be a sequence of nonnegative real numbers such that $\lim_{q\rightarrow\infty}\theta_{h,q}=\theta_{h}$.
+For $h\in\mathbf{N}$, let $\theta_{h}\geq 0$ and let $(\theta_{h,q})_{q=1}^{\infty}$ be a decreasing sequence of nonnegative real numbers such that $\lim_{q\rightarrow\infty}\theta_{h,q}=\theta_{h}$.
 Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of compact subsets of $G$ such that $\mu(hA_{q})=\theta_{h,q}$ for all $q\in\mathbf{N}$.
@@
 and so
-$\mu(hA)=\mu\left(\bigcap_{q=1}^{\infty}hA_{q}\right)=\lim_{q\rightarrow\infty}\mu(hA_{q})=\lim_{q\rightarrow\infty}\mu(\theta_{h,q})=\theta_{h}.$
+$\mu(hA)=\mu\left(\bigcap_{q=1}^{\infty}hA_{q}\right)=\lim_{q\rightarrow\infty}\mu(hA_{q})=\lim_{q\rightarrow\infty}\theta_{h,q}=\theta_{h}.$
 This completes the proof. ∎
-4. Semigroup problems
-Problem 5.
+4. Additional problems
+Problem 3.
 Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
 Describe the set $\{h\in\mathbf{N}:hA=\bigcap_{q=1}^{\infty}hA_{q}\}$.
-Problem 6.
+What sets of integers arise in this way?
+Problem 4.
 Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
@@
 When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h-1)A=\bigcap_{q=1}^{\infty}(h-1)A_{q}$?
-I thank Diego Marques for the observation that led to Theorem 2.
+Problem 5.
+For what sets $A$ of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
+Acknowledgement. I thank Diego Marques for helpful comments that suggested Theorem 6.